What is the explicit formula of 5, 11, 23, 47?
Answer by josgarithmetic(39616)

5+6=11
11+6*2=11+12=23
23+6*4=23+24=47

Predict that the next term would be 47+6*6, but this is not certain; and the four terms given are not yet enough for an unambiguous pattern.